Circular No. 8539

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET 9P/TEMPEL H. U. Kaeufl, European Southern Observatory (ESO); T. Bonev, Insti-

tute of Astronomy, Bulgarian Academy of Sciences and visiting astronomer at ESO; H. Boenhardt, Max Planck Institute for Solar System Research; Y. R. Fernandez, Institute of Astronomy, University of Hawaii; and C. Lisse, University of Maryland, report that mid-infrared images of comet 9P were obtained at the ESO 3.6-m telescope (+ TIMMI2; spatially resolved at a sub-arcsec level), and Bessel R-band images were obtained with the ESO 3.5-m New Technology Telescope (+ EMMI). Fluxes derived from midinfrared images at 8.9, 9.8, 10.4, and 11.9 μm, respectively, in a circular aperture of radius 3′′ in units of 10^{−15} W m^{−2} μm^{−1}: Feb. 23, 6.0, 5.2, 6.6, 6.0; Mar. 25, 16.4, 20.0, 20.0, 22.6. R-band (Bessel) fluxes measured in a circular aperture of radius 3′′, also in units of 10^{−15} W m^{−2} μm^{−1}: Feb. 23, 13; Mar. 25, 32. A prominent straight jet is seen in the R-band images at p.a. 220^{o} and 215^{o} on Feb. 23 and Mar. 25, respectively.

SUPERNOVA 2005ch Further to IAUC 8537, E. S. Rykoff reports the discovery of an ap-

parent supernova in unfiltered CCD images taken on June 8.83 (at mag ∼ 15.8) and 9.56 UT (mag ∼ 15.5) with the 0.45-m ROTSE-IIIc telescope at the HESS site in Namibia and the ROTSE-IIIa telescope at Siding Spring Observatory. SN 2005ch is located at α = 14^{h}22^{m}06^{s}.77, δ = +1^{o}59′35′′.7 (equinox 2000.0), which is 0′′.9 north and 3′′.1 east of the core of the apparent host galaxy (whose Sloan Digital Sky Survey magnitude is given as g′ = 20.0). Nothing is visible at this location on ROTSE-IIIc images taken on May 27.86 (limiting mag ∼ 17.5).

VARIABLE STAR IN NORMA Further to IAUC 8505, G. Pojmanski announces the discovery on ASAS

images of a variable star located at α = 16^{h}00^{m}48^{s}, δ = −48^{o}46′.2 (equinox 2000.0). ASAS V magnitudes for the variable: June 4.316 UT, [14 (limiting mag presumed from earlier publications); 9.085, 12.57; 9.235, 12.66; 9.288, 12.82; 9.296, 12.67; 9.321, 12.81. A USNO-A1.0 star of red mag 18.4 is located at position end figures 48^{s}.233, 07′′.21.

COMET P/2005 JY_{126} (CATALINA) Corrigendum. On IAUC 8537, line 8, for p.a. 65^{o}. read p.a. 245^{o}.

2005 June 9  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2005 CBAT  Daniel W. E. Green